# Geometry-aware Temporal Aggregation Network for Monocular 3D Lane Detection

### v1

@@ 3.3 Temporal Instance-aware Query Generation
Successive frames provide different perspectives of the driving scene. Each frame contains unique contextual cues and dynamic changes within the scene. In addition to the geometric consistency between successive frames, we also aim to leverage the information variety between them, revealing rich instance information to alleviate the difficulties in maintaining lane integrity. Hence, we propose the Temporal Instance-aware Query Generation (TIQG), as shown in figure 3. There are three main steps in TIQG including initial query generation, synthetic future query generation and temporal query aggregation.
@@
Table 1: Quantitative Comparison of Monocular 3D Lane Detection Solutions on the OpenLane Dataset. $\downarrow$ and $\uparrow$ indicate that lower and higher values, respectively, signify superior performance. Bold numbers represent the best results, while underlined numbers indicate the second best. GTA-Net obtains the competitive performance on F1-score, lane category accuracy, x/z errors for both near and far range.
      Category       X error (m) $\downarrow$       Z error (m) $\downarrow$
      Methods         F1 (%) $\uparrow$       Accuracy (%) $\uparrow$       near        far       near        far
      3DLaneNet [10]       44.1       -       0.479       0.572       0.367       0.443
      GenLaneNet [13]       32.3       -       0.593       0.494       0.140       0.195
      Cond-IPM       36.6       -       0.563       1.080       0.421       0.892
      Persformer [5]       50.5       89.5       0.319       0.325       0.112       0.141
      CurveFormer [2]       50.5       -       0.340       0.772       0.207       0.651
      BEV-LaneDet [48]       58.4       -       0.309       0.659       0.244       0.631
      Anchor3DLane [17]       53.1       90.0       0.300       0.311       0.103       0.139
      DecoupleLane [15]       51.2       -       -       -       -       -
      CurveFormer++ [3]       52.5       87.8       0.333       0.805       0.186       0.687
      LaneCMK [57]       55.8       89.2       0.310       0.303       0.083       0.123
      GroupLane [26]       60.2       91.6       0.371       0.476       0.220       0.357
      CaliFree3DLane [12]       56.4       -       0.319       0.767       0.245       0.706
      CaliFree3DLane [12]       57.0       -       0.306       0.761       0.227       0.683
      3DLaneFormer [7]       55.0       89.9       0.283       0.305       0.106       0.138
      Ours       62.4       92.8       0.225       0.254       0.078       0.110
Initial query generation. The initial query generation process involves generating temporal queries for both frame $t$ and frame $t-n$. The queries consist of two-level embeddings: instance-level embeddings and point-level embeddings. To derive instance-level embedding, cascaded 2D convolutions are employed to process image features $F_{t}$ and $F_{t-n}$ as follows:
$Q_{t}^{l}=\text{Conv}(F_{t}),\quad Q_{t-n}^{l}=\text{Conv}(F_{t-n}),\$ (14)
where $Q_{t}^{l}$ and $Q_{t-n}^{l}$ denote the generated instance-level embeddings of frame $t$ and frame $t-n$, respectively. For point-level embedding, we define two learnable parameters $Q_{t}^{p}$ and $Q_{t-n}^{p}$. Subsequently, these corresponding two-level embeddings are merged using the following formulation:
$Q_{t}=Q_{t}^{l}\oplus Q_{t}^{p},\quad Q_{t-n}=Q_{t-n}^{l}\oplus Q_{t-n}^{p},\$ (15)
where $\oplus$ represents the broadcasting summation.
@@
     Methods      All      Up &      Curve      Extreme      Night      Intersection      Merge
     Down      Weather      & Split
     3DLaneNet [10]      44.1      40.8      46.5      47.5      41.5      32.1      41.7
     GenLaneNet [13]      32.3      25.4      33.5      28.1      18.7      21.4      31.0
     Persformer [5]      50.5      42.4      55.6      48.6      46.6      40.0      50.7
     CurveFormer [2]      50.5      45.2      56.6      49.7      49.1      42.9      45.4
     BEV-LaneDet [48]      58.4      48.7      63.1      53.4      53.4      50.3      53.7
     Anchor3DLane [17]      54.3      47.2      58.0      52.7      48.7      45.8      51.7
     DecoupleLane [15]      51.2      43.5      57.3      -      48.9      43.5      -
     LATR [35]      61.9      55.2      68.2      57.1      55.4      52.3      61.5
     CurveFormer++ [3]      52.7      48.3      59.4      50.6      48.4      45.0      48.1
     LaneCMK [57]      55.8      47.3      58.6      53.2      48.0      42.2      51.7
     CaliFree3DLane [12]      56.4      48.7      60.4      54.7      51.9      49.5      51.8
     CaliFree3DLane [12]      57.0      48.9      62.3      54.8      52.0      50.2      52.1
     3DLaneFormer [7]      55.0      49.0      58.3      54.5      50.4      46.3      52.3
     Anchor3DLane++ [18]      59.4      49.9      66.1      55.5      52.5      50.3      57.4
     Ours      62.4      55.5      68.3      57.5      56.8      52.7      61.8
@@
$Q_{p}^{u}=\text{CA}(Q_{p},Q_{t}),\quad Q_{t-1}^{u}=\text{CA}(Q_{t-1},Q_{t}),\$ (17)
where $Q_{p}^{u}$, $Q_{t-1}^{u}$ and $\text{CA}(\cdot)$ denote the updated synthetic future queries, the updated queries of frame $t-1$ and the transformation of cross-attention. Subsequently, these lane queries are fed into cascaded convolutions to obtain instance-aware queries $Q$:
$Q=\text{Conv}(\text{Concat}(Q_{t},Q_{t-1},Q_{p})),\$ (18)
@@ 3.4 Loss Function
With the matched ground truth 3D lanes corresponding to the output 3D lane queries, we compute the loss for each respective pair. To be specific, L1 loss is employed to regress the $x$-coordinates and $z$-coordinates. Binary cross-entropy loss is used to enforce the visibility constraint of the output 3D lanes. Focal loss [28] is utilized to optimize the predicted lane category. In addition, a 2D segmentation loss is also included as an auxiliary loss. Hence, the final objective function of our GTA-Net can be formulated as follows:
@@ 4 Experiments
This section provides a detailed overview of our experimental setup. We then evaluate the performance of our GTA-Net through a series of comparative analyses against state-of-the-art methods. Additionally, we present several ablation studies to gain a deeper understanding of the effectiveness of our proposed approach.
@@ 4.1 Experimental Setup
Dataset and Evaluation metrics. In the experiments, we use the OpenLane dataset [5], derived from the Waymo dataset [42], as a comprehensive benchmark for 3D lane detection. Comprising 1,000 segments and 200,000 frames, this dataset captures diverse environmental conditions such as varying weather, terrain, and lighting. The images in the OpenLane dataset have a resolution of 1280$\times$1920 pixels and include 880,000 lane annotations spanning 14 categories. The OpenLane dataset provides a diverse and challenging set for evaluating 3D lane detection algorithms. We employ the official evaluation metrics to assess the effectiveness of our model on the OpenLane dataset. Specifically, we report the F1 score, lane category accuracy, and x/z errors in both the near and far ranges.
Implementation details. In our GTA-Net, we use ResNet-50 [16] as the 2D backbone network for extracting image features. Before feature extraction, all input images are resized to a resolution of 720$\times$960. The experiments are conducted on the PyTorch platform using a workstation equipped with eight Tesla A100 GPUs, each with 40 GB of memory. For optimization, we employ the AdamW optimizer with a weight decay of 0.01. The initial learning rate is set to $2e^{-4}$, and a cosine annealing scheduler is applied to gradually adjust the learning rate. Our model is trained for 24 epochs, with a batch size of 64. Additionally, we set $w_{x}=2$, $w_{z}=10$, $w_{cate}=10$ and $w_{seg}=5$.
Compared methods. We compare our GTA-Net with the following monocular 3D detection methods: 3DLaneNet [10], Gen-LaneNet [13], PersFormer [5], Cond-IPM, CurveFormer [2], DecoupleLane [15], Anchor3DLane [17], BEV-LaneDet [48], LATR [35], GroupLane [26], LaneCMK [57], CaliFree3DLane [12], LaneCPP [38], 3DLaneFormer [7] and CurveFormer++ [3].
@@
Table 3: Ablation Study of the Proposed Novel Modules Including TGEM and TIQG. We have designed a series of variants to systematically evaluate the effectiveness of each individual module.
      Category       X error (m) $\downarrow$       Z error (m) $\downarrow$
      Methods         F1 (%) $\uparrow$       Accuracy (%) $\uparrow$       near        far       near        far
      w/o. TGEM       62.1       92.1       0.229       0.257       0.081       0.112
      w/o. TIQG       62.1       91.7       0.229       0.264       0.080       0.114
      Ours       62.4       92.8       0.225       0.254       0.078       0.110
@@ 4.2 Main Results
Quantitative results. We present the experimental results of our method on the OpenLane dataset in Table 1. As shown in the results, our proposed GTA-Net outperforms existing state-of-the-art methods for 3D lane detection. Specifically, GTA-Net achieves the highest F1 score and lane category accuracy, underscoring its superior capability in accurately detecting 3D lanes while minimizing both missed detections and false positives. Moreover, when evaluating the localization of the detected 3D lanes, our method achieves competitive x/z errors under both far and near range in comparison to other monocular 3D lane detection approaches, further validating its effectiveness for perceiving both close and distant lanes.
To further validate the effectiveness of the proposed GTA-Net, we present results across a range of challenging scenarios. Table 2 reports the 3D lane detection performance on the OpenLane dataset under different conditions, measured by the F1 score. Notably, our method outperforms existing approaches, achieving the highest F1 score across all scenarios. This demonstrates the robustness of our model in handling diverse driving environments, underscoring its reliability and efficacy in accurately detecting 3D lanes under varying real-world conditions.
Visual results. To rigorously assess the performance of GTA-Net, we present the visualization results on the OpenLane dataset. As depicted in figure 5, our model demonstrates the capability to accurately detect 3D lanes in the input images and to precisely estimate their corresponding positions in 3D space. These results illustrate the excellent performance of our model in accurately identifying and localizing lanes in 3D space.
@@ 4.3 Ablation Studies
Effect of TGEM. We first evaluate the contribution of the proposed TGEM module. As presented in Table 3, we assess the performance of GTA-Net without the TGEM module, denoted as w/o TGEM. The results demonstrate a marked degradation in performance, highlighting the critical role of TGEM in promoting the ability to capture and leverage geometric cues from the scene.
Effect of TIQG. We next evaluate the contribution of the TIQG module. As presented in Table 3, the variant labeled w/o TIQG corresponds to a version of GTA-Net where the TIQG module is omitted. The results reveal a significant degradation in performance across all metrics when TIQG is excluded, thereby suggesting its critical role in effectively leveraging temporal cues. This demonstrates the importance of TIQG in improving the accuracy of 3D lane detection from monocular inputs by harnessing temporal instance information to enhance the integrity of lanes.
@@ 5 Conclusion
In this paper, we tackle two key challenges in monocular 3D lane detection: the inaccurate geometric information of predicted 3D lanes and the difficulty in detecting remote lanes. To address these issues, we propose a novel Geometry-aware Temporal Aggregation Network (GTA-Net) for monocular 3D lane detection. We introduce the Temporal Geometry Enhancement Module (TGEM), which leverages geometric consistency across successive frames to enhance depth perception and improve scene understanding. Additionally, we present the Temporal Instance-aware Query Generation (TIQG) module, which strategically incorporates temporal cues into the query generation process, enabling a more comprehensive exploration of instance-aware information across frames. Experimental results demonstrate that the proposed GTA-Net outperforms existing monocular 3D lane detection methods.

### v3

@@ 3.3 Temporal Instance-aware Query Generation
While the TGEM module provides rich geometric context (what the scene looks like), it does not explicitly tell us where the lanes are. Accurately localizing all lanes, especially those that are distant or partially occluded, requires instance-level priors. Successive frames offer a wealth of such information: a lane barely visible now might have been clear seconds ago. To harness this rich temporal context, we propose the Temporal Instance-aware Query Generation (TIQG) module, as illustrated in Figure 3. The goal of TIQG is to generate a comprehensive and highly accurate set of lane queries by strategically aggregating information from the past, present, and a synthesized future. This process unfolds in three key stages: initial query generation from monocular frames, synthetic future query generation for distant lanes, and finally, temporal query aggregation.
Initial query generation. The initial query generation process involves generating temporal queries for both frame $t$ and frame tnt-n. The queries consist of two-level embeddings: instance-level embeddings and point-level embeddings. To derive instance-level embedding, cascaded 2D convolutions are employed to process image features $F_{t}$ and $F_{t-n}$ as follows:
$Q_{t}^{l}=\text{Conv}(F_{t}),\quad Q_{t-n}^{l}=\text{Conv}(F_{t-n}),\\$ (14)
where $Q_{t}^{l}$ and $Q_{t-n}^{l}$ denote the generated instance-level embeddings of frame $t$ and frame tnt-n, respectively. For point-level embedding, we define two learnable parameters $Q_{t}^{p}$ and $Q_{t-n}^{p}$. Subsequently, these corresponding two-level embeddings are merged using the following formulation:
$Q_{t}=Q_{t}^{l}\oplus Q_{t}^{p},\quad Q_{t-n}=Q_{t-n}^{l}\oplus Q_{t-n}^{p},\\$ (15)
where $\oplus$ represents the broadcasting summation.
@@
Table 1: Quantitative Comparison of Monocular 3D Lane Detection Solutions on the OpenLane Dataset. $\downarrow$ and $\uparrow$ indicate that lower and higher values, respectively, signify superior performance. Bold numbers represent the best results. GTA-Net obtains the competitive performance on F1-score, lane category accuracy, x/z errors for both near and far range.
   Category      X error (m) $\downarrow$    Z error (m) $\downarrow$
   Methods      Year      F1 (%) $\uparrow$    Accuracy (%)   $\uparrow$    near     far    near     far
   3DLaneNet [garnett20193d]    2019    44.1    -    0.479    0.572    0.367    0.443
   GenLaneNet [guo2020gen]    2020    32.3    -    0.593    0.494    0.140    0.195
   Cond-IPM    -    36.6    -    0.563    1.080    0.421    0.892
   Persformer [chen2022persformer]    2022    50.5    89.5    0.319    0.325    0.112    0.141
   CurveFormer [bai2023curveformer]    2023    50.5    -    0.340    0.772    0.207    0.651
   BEV-LaneDet [wang2022bev]    2022    58.4    -    0.309    0.659    0.244    0.631
   Anchor3DLane [huang2023anchor3dlane]    2023    53.1    90.0    0.300    0.311    0.103    0.139
   DecoupleLane [han2023decoupling]    2023    51.2    -    -    -    -    -
   LaneCMK [zhao2024lanecmkt]    2024    55.8    89.2    0.310    0.303    0.083    0.123
   GroupLane [li2024grouplane]    2024    60.2    91.6    0.371    0.476    0.220    0.357
   CaliFree3DLane [guo2024califree3dlane]    2024    56.4    -    0.319    0.767    0.245    0.706
   CaliFree3DLane [guo2024califree3dlane]    2024    57.0    -    0.306    0.761    0.227    0.683
   3DLaneFormer [dong20243dlaneformer]    2024    55.0    89.9    0.283    0.305    0.106    0.138
   DB3D-L [liu2025db3d]    2025    55.2    -    0.254    0.682    0.202    0.622
   Freq-3DLane [song2025freq]    2025    59.7    -    0.265    0.665    0.196    0.608
   CurveFormer++ [bai2025curveformer++]    2025    52.5    87.8    0.333    0.805    0.186    0.687
   Ours    -    62.4    92.8    0.225    0.254    0.078    0.110
@@
$Q_{p}^{u}=\text{CA}(Q_{p},Q_{t}),\quad Q_{t-1}^{u}=\text{CA}(Q_{t-1},Q_{t}),\\$ (17)
where $Q_{p}^{u}$, $Q_{t-1}^{u}$ and $\text{CA}(\cdot)$ denote the updated synthetic future queries, the updated queries of frame t1t-1 and the transformation of cross-attention. Subsequently, these lane queries are fed into cascaded convolutions to obtain instance-aware queries $Q$:
$Q=\text{Conv}(\text{Concat}(Q_{t},Q_{t-1},Q_{p})),\\$ (18)
@@ 3.4 Loss Function
With the matched ground truth 3D lanes corresponding to the output 3D lane queries, we compute the loss for each respective pair. To be specific, L1 loss is employed to regress the $x$-coordinates and $z$-coordinates. Binary cross-entropy loss is used to enforce the visibility constraint of the output 3D lanes. Focal loss [lin2017focal] is utilized to optimize the predicted lane category. In addition, a 2D segmentation loss is also included as an auxiliary loss. Hence, the final objective function of our GTA-Net can be formulated by:
@@
Methods Year All Up & Curve Extreme Night Intersection Merge
Down Weather & Split
3DLaneNet [garnett20193d] 2019 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [guo2020gen] 2020 32.3 25.4 33.5 28.1 18.7 21.4 31.0
Persformer [chen2022persformer] 2022 50.5 42.4 55.6 48.6 46.6 40.0 50.7
CurveFormer [bai2023curveformer] 2023 50.5 45.2 56.6 49.7 49.1 42.9 45.4
BEV-LaneDet [wang2022bev] 2022 58.4 48.7 63.1 53.4 53.4 50.3 53.7
Anchor3DLane [huang2023anchor3dlane] 2023 54.3 47.2 58.0 52.7 48.7 45.8 51.7
DecoupleLane [han2023decoupling] 2023 51.2 43.5 57.3 - 48.9 43.5 -
LATR [luo2023latr] 3 2023 61.9 55.2 68.2 57.1 55.4 52.3 61.5
LaneCMK [zhao2024lanecmkt] 2024 55.8 47.3 58.6 53.2 48.0 42.2 51.7
CaliFree3DLane [guo2024califree3dlane] 2024 56.4 48.7 60.4 54.7 51.9 49.5 51.8
CaliFree3DLane [guo2024califree3dlane] 2024 57.0 48.9 62.3 54.8 52.0 50.2 52.1
3DLaneFormer [dong20243dlaneformer] 2024 55.0 49.0 58.3 54.5 50.4 46.3 52.3
Freq-3DLane [song2025freq] 2025 59.7 52.0 65.5 56.5 54.5 51.8 56.4
CurveFormer++ [bai2025curveformer++] 5 2025 52.7 48.3 59.4 50.6 48.4 45.0 48.1
Ours - 62.4 55.5 68.3 57.5 56.8 52.7 61.8
@@ 4 Experiments
This section provides a detailed overview of our experimental setup. We then evaluate the performance of our GTA-Net through a series of comparative analyses against state-of-the-art methods. Additionally, we present several ablation studies to gain a deeper understanding of the effectiveness of our proposed approach. Through these experiments, we aim to provide a thorough assessment of the capabilities and contributions of our model.
@@ 4.1 Experimental Setup
[…]
Compared methods. We compare our GTA-Net with the following monocular 3D detection methods: 3DLaneNet [garnett20193d], Gen-LaneNet [guo2020gen], PersFormer [chen2022persformer], Cond-IPM, CurveFormer [bai2023curveformer], DecoupleLane [han2023decoupling], Anchor3DLane [huang2023anchor3dlane], BEV-LaneDet [wang2022bev], LATR [luo2023latr], GroupLane [li2024grouplane], LaneCMK [zhao2024lanecmkt], CaliFree3DLane [guo2024califree3dlane], LaneCPP [pittner2024lanecpp], 3DLaneFormer [dong20243dlaneformer], DB3D-L [liu2025db3d], Freq-3DLane [song2025freq] and CurveFormer++ [bai2025curveformer++].
@@ 4.2 Main Results
Quantitative results. We present a comprehensive quantitative evaluation of our proposed GTA-Net on the challenging OpenLane dataset, with detailed results summarized in Table 1. The experimental findings unequivocally demonstrate that GTA-Net achieves superior performance, setting a new SoTA benchmark for monocular 3D lane detection.
Specifically, GTA-Net excels across key evaluation metrics. It attains the highest F1 score, indicative of its robust ability to precisely identify 3D lanes while effectively balancing recall and precision to minimize both missed detections and false positives. Furthermore, our method exhibits superior lane category accuracy, underscoring its enhanced capability in semantically understanding and classifying diverse lane types. Beyond detection, GTA-Net also demonstrates exceptional localization precision. It consistently yields the smallest $x$ and $z$ errors across both near and far perception ranges, a critical validation of its effectiveness in accurately estimating the spatial coordinates of 3D lanes, irrespective of their distance from the ego-vehicle. These results collectively affirm the robustness and practical utility of GTA-Net in complex scenarios.
To further validate the effectiveness of the proposed GTA-Net, we present results across a range of challenging scenarios. Table 2 reports the 3D lane detection performance under different conditions, measured by the F1 score. Notably, our method outperforms existing approaches, achieving the highest F1 score across all scenarios. This demonstrates the robustness of our model in handling diverse driving environments, underscoring its reliability and efficacy in accurately detecting 3D lanes under varying real-world conditions.
Visual Results. Beyond quantitative metrics, a rigorous qualitative assessment is crucial for a comprehensive understanding of GTA-Net’s performance. To this end, we present a series of visualization results on the challenging OpenLane dataset, offering an intuitive insight into our model’s capabilities. As vividly depicted in Figure 5, GTA-Net consistently demonstrates an exceptional ability to accurately detect 3D lanes directly within the complex input images. More importantly, it precisely estimates their corresponding positions in 3D space, even under varying scene conditions. These visual outcomes not only corroborate the superior quantitative performance but also powerfully illustrate the model’s robustness and accuracy in real-world scenarios, affirming its excellent capacity for both identifying and accurately localizing lanes in 3D space.
Table 3: Ablation study of the proposed modules. We evaluate the effect of TGEM and TIQG by ablating them.
Methods F1 (%) X error (m) Z error (m)
near far near far
Baseline 61.7 0.235 0.273 0.085 0.120
w/o. TGEM 62.1 0.229 0.257 0.081 0.112
w/o. TIQG 62.1 0.229 0.264 0.080 0.114
w/o. Cost Volume 62.2 0.228 0.256 0.080 0.112
Ours 62.4 0.225 0.254 0.078 0.110
@@ 4.3 Ablation Studies
Effect of TGEM. To systematically dissect the contributions of individual components within our proposed GTA-Net framework, we first conduct an ablation study focusing on the Temporal Geometry Enhancement Module (TGEM). As detailed in Table 3, we evaluate the performance of a modified GTA-Net variant, denoted as w/o. TGEM, where the TGEM module is deliberately omitted. The comparative results reveal a marked and consistent degradation in performance across all evaluated metrics when TGEM is excluded. This significant decline unequivocally highlights the critical and indispensable role of TGEM. It actively promotes the network’s ability to accurately capture and robustly leverage intricate geometric cues from the temporal evolution of the scene, which are crucial for resolving ambiguities inherent in monocular 3D lane detection.
Effect of TIQG. We next evaluate the contribution of the TIQG module. As presented in Table 3, the variant labeled w/o. TIQG corresponds to a version of GTA-Net where the TIQG module is omitted. The results reveal a significant degradation in performance across all metrics when TIQG is excluded, thereby suggesting its critical role in effectively leveraging temporal cues. This demonstrates the importance of TIQG in improving the accuracy of 3D lane detection from monocular inputs by harnessing temporal instance information to enhance the integrity of lanes.
Effect of Cost Volume.
To investigate the contribution of the cost volume construction within TGEM, we conduct an ablation study where this component is omitted. In this variant, referred to as w/o. Cost Volume, the geometric feature learning directly processes the concatenated raw 2D features without an explicit cost volume generation step. As shown in Table 3, the removal of the cost volume leads to a degradation in performance across all metrics. This suggests that while cost volume plays a role in capturing precise geometric relationships between frames. The cost volume provides a fine-grained inter-frame geometric correspondence that aids in depth-aware feature generation, and its absence impacts the overall accuracy.
@@ 5 Conclusion
In this paper, we tackle two key challenges in monocular 3D lane detection: the inaccurate geometric information of predicted 3D lanes and the difficulty in detecting remote lanes. To address these issues, we propose a novel Geometry-aware Temporal Aggregation Network (GTA-Net) by harnessing temporal evolution for monocular 3D lane detection. We introduce the Temporal Geometry Enhancement Module (TGEM), which leverages geometric consistency across successive frames to enhance depth perception and improve geometric understanding of the scene. Additionally, we present the Temporal Instance-aware Query Generation (TIQG) module, which strategically incorporates temporal context into the query generation process, enabling a more comprehensive exploration of instance-aware information across frames. Through extensive and rigorous experimentation, the results demonstrate that the proposed GTA-Net outperforms existing methods.